\begin{thebibliography}{10}
\providecommand{\url}[1]{#1}
\csname url@rmstyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand\BIBentrySTDinterwordspacing{\spaceskip=0pt\relax}
\providecommand\BIBentryALTinterwordstretchfactor{4}
\providecommand\BIBentryALTinterwordspacing{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand\BIBforeignlanguage[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}

\bibitem{Nechaev82}
A.~A. Nechaev, ``Trace-function in {G}alois ring and noise-stable codes,'' in
  \emph{Proc. V All-Union Symp. on Theory of Rings, Alg. and Mod.},
  Novosibirsk, Russia, 1982, p.~97, in Russian.

\bibitem{Nechaev:Kerdock_cyclic_form}
A.~A. Nechaev, ``Kerdock code in a cyclic form,'' \emph{Discrete Math. Appl.}, vol.~1, no.~4, pp. 365--384, 1991. Translated from {\it
  Diskretnaya Matematika}, \href{http://mi.mathnet.ru/rus/dm/v1/i4/p123}{1(4): 123--139, 1989}.

\bibitem{HammonsOth:Z4_linearity}
A.~R.~J. Hammons, P.~V. Kumar, A.~R. Calderbank, N.~J.~A. Sloane, and
  P.~Sol\'e, ``The {$Z_4$}-linearity of {K}erdock, {P}reparata, {G}oethals, and
  related codes,'' \emph{IEEE Trans. Inform. Theory}, vol.~40, no.~2, pp.
  301--319, 1994. DOI:
  \href{http://dx.doi.org/10.1109/18.312154}{10.1109/18.312154}

\bibitem{Carlet:Z2k-linear}
C.~Carlet, ``{$Z_{2^k}$}-linear codes,'' \emph{IEEE Trans. Inform. Theory},
  vol.~44, no.~4, pp. 1543--1547, 1998.
  DOI: \href{http://dx.doi.org/10.1109/18.681328}{10.1109/18.681328}

\bibitem{HonNech}
T.~Honold and A.~A. Nechaev, ``Fully weighted modules and representations of
  codes,'' \emph{Probl. Inform. Transm.}, vol.~35, no.~3, pp. 205--223, 1999.
Translated from \emph{Probl. Peredachi Inf.},
\href{http://mi.mathnet.ru/rus/ppi/v35/i3/p18}{35(3): 18--39, 1999}.

\bibitem{ConstHeise}
I.~Constantinescu and W.~Heise, ``A metric for codes over residue class rings
  of integers,'' \emph{Probl. Inform. Transm.}, vol.~33, no.~3, pp. 208--213, 1997.
Translated from \emph{Probl. Peredachi Inf.},
\href{http://mi.mathnet.ru/rus/ppi/v33/i3/p22}{33(3): 22--28, 1997}.

\bibitem{Zin1976:GCC}
V.~A. Zinoviev, ``Generalized concatenated codes,'' \emph{Probl. Inform.
  Transm.}, vol.~12, no.~1, pp. 2--9, 1976. Translated from \emph{Probl.
  Peredachi Inf.},
\href{http://mi.mathnet.ru/rus/ppi/v12/i1/p5}{12(1): 5-15, 1976}.

\bibitem{RifPuj:1997}
J.~Rifa and J.~Pujol, ``Translation-invariant propelinear codes,'' \emph{IEEE
  Trans. Inform. Theory}, vol.~43, no.~2, pp. 590--598, 1997.
  DOI: \href{http://dx.doi.org/10.1109/18.556115}{10.1109/18.556115}

\bibitem{BorRif:1999}
J.~Borges and J.~Rifa, ``A characterization of $1$-perfect additive codes,''
  \emph{IEEE Trans. Inform. Theory}, vol.~45, no.~8, pp. 1688--1697, 1999.
  DOI: \href{http://dx.doi.org/10.1109/18.771247}{10.1109/18.771247}

\bibitem{Kro:2000:Z4_Perf}
D.~S. Krotov, ``{$Z_4$}-linear perfect codes,'' \emph{Diskr. Analiz i Issled.
  Operatsii, Ser.1},
\href{http://mi.mathnet.ru/rus/da/v7/s1/i4/p78}{vol.~7, no.~4, pp. 78--90, 2000}, in Russian. English translation: \href{http://arXiv.org/abs/0710.0198}{arXiv:0710.0198}

\bibitem{Kro:2001:Z4_Had_Perf}
D.~S. Krotov, ``{$Z_4$}-linear {H}adamard and extended perfect codes,''
  \emph{Electronic Notes in Discrete Mathematics}, vol.~6, pp. 107--112, 2001.
 DOI:
   \href{http://dx.doi.org/10.1016/S1571-0653(04)00161-1}{10.1016/S1571-0653(04)00161-1} \href{http://arXiv.org/abs/0710.0199}{arXiv:0710.0199}

\bibitem{PheRif:2002}
K.~T. Phelps and J.~Rifa, ``On binary $1$-perfect additive codes: Some
  structural properties,'' \emph{IEEE Trans. Inform. Theory}, vol.~48, no.~9,
  pp. 2587--2592, 2002. DOI:
  \href{http://dx.doi.org/10.1109/TIT.2002.801474}{10.1109/TIT.2002.801474}

\bibitem{BorPheRif:2003}
J.~Borges, K.~T. Phelps, and J.~Rifa, ``The rank and kernel of extended
  $1$-perfect {$Z_4$}-linear and additive non-{$Z_4$}-linear codes,''
  \emph{IEEE Trans. Inform. Theory}, vol.~49, no.~8, pp. 2028--2034, 2003.
  DOI: \href{http://dx.doi.org/10.1109/TIT.2003.814490}{10.1109/TIT.2003.814490}

\bibitem{PheRifVil:2004ACCT:additive_hadamard}
K.~T. Phelps, J.~Rifa, and M.~Villanueva, ``Rank and kernel of additive
  ({$Z_4$}-linear and non-{$Z_4$}-linear) {H}adamard codes,'' in \emph{Proc.
  {N}inth Int. Workshop on Algebraic and Combinatorial Coding Theory
  {ACCT'2004}}, Kranevo, Bulgaria, June 2004, pp. 327--332.

\bibitem{PheRifVil:Hadamard-rank-kernel}
K.~T. Phelps, J.~Rifa, and M.~Villanueva, ``On the additive ({$Z_4$}-linear and non-{$Z_4$}-linear) {H}adamard
  codes. {R}ank and kernel,'' \emph{IEEE Trans. Inform. Theory}, vol.~52, no.~1, pp. 316--319, 2006.
  DOI:
  \href{http://dx.doi.org/10.1109/TIT.2005.860452}{10.1109/TIT.2005.860452}

\bibitem{GuptaOth2005}
M.~K. Gupta, M.~C. Bhandari, and A.~K. Lal, ``On linear codes over
  {$Z_{2^s}$},'' \emph{Des. Codes Cryptography}, vol.~36, no.~3, pp.
  227--244, 2005. DOI:
  \href{http://dx.doi.org/10.1007/s10623-004-1717-1}{10.1007/s10623-004-1717-1}

\bibitem{Sol:2005transitive}
F.~I. Solov'eva, ``On the construction of transitive codes,'' \emph{Probl.
  Inform. Transm.}, vol.~41, no.~3, pp. 204--211, 2005.
  DOI: \href{http://dx.doi.org/10.1007/s11122-005-0025-3}{10.1007/s11122-005-0025-3}
Translated from \emph{Probl. Peredachi Inf.},
\href{http://mi.mathnet.ru/rus/ppi/v41/i3/p23}{41(3): 23--31, 2005}.

\bibitem{MWS}
F.~J. MacWilliams and N.~J.~A. Sloane, \emph{The Theory of Error-Correcting
  Codes}.\hskip 1em plus 0.5em minus 0.4em\relax Amsterdam, Netherlands: North
  Holland, 1977.

\bibitem{AhlAydKha}
R.~Ahlswede, H.~K. Aydinian, and L.~K. Khachatrian, ``On perfect codes and
  related concepts,'' \emph{Des. Codes Cryptography}, vol.~22, no.~3,
  pp. 221--237, 2001. DOI:  \href{http://dx.doi.org/10.1023/A:1008394205999}{10.1023/A:1008394205999}

\end{thebibliography}

\documentclass[letterpaper,journal,onecolumn,oneside,final,12pt]{IEEEtran}
\usepackage{amssymb} 
\usepackage{color}
\definecolor{dkgreen}{rgb}{0,.5,0}
\definecolor{dkred}{rgb}{.5,0,0}
\definecolor{dkcyan}{rgb}{0,.5,.5}
\definecolor{dkmgnt}{rgb}{.5,0,.5}
\usepackage[dvips,colorlinks,linkcolor=dkred,pagecolor=red,filecolor=dkcyan,citecolor=dkgreen,urlcolor=dkmgnt%
,bookmarksnumbered=true%
,pdftitle=On\ $Z_{2^k}$-Dual\ Binary\ Codes%
,pdfsubject=Coding\ Theory%
,pdfauthor=Denis\ S.\ Krotov%
,pdfkeywords=Gray\ map;\ Hadamard\ codes;\ MacWilliams\ identity;\ perfect\ codes;\ $Z_{2^k}$-linearity%
,hypertexnames=false%
]{hyperref} 

\providecommand\hyperref[2][]{#2} 
\providecommand\triangleq{{\stackrel{\scriptscriptstyle\triangle}{{}={}}}} 

\renewcommand{\QED}{\QEDclosed} 
\newcommand{\proofr}{\begin{proof}} 
\newcommand{\proofrem}[1]{\par\indent\indent \emph{#1:}} 
\newcommand{\proofend}{\end{proof}}  
\newcommand{\proofremend}{\hfill\QED}  

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{pro}[theorem]{Proposition}
\newtheorem{corol}[theorem]{Corollary}
\newtheorem{note}[theorem]{Remark}
\newtheorem{examp}[theorem]{Example}

\newcommand{\bpro}{\begin{pro}}
\newcommand{\bprorem}[1]{\begin{pro}[#1]}
\newcommand{\epro}{\end{pro}}

\newcommand{\bcor}{\begin{corol}}
\newcommand{\ecor}{\end{corol}}

\newcommand{\bnote}{\begin{note}}
\newcommand{\enote}{\end{note}}

\newcommand{\btheorem}{\begin{theorem}}
\newcommand{\etheorem}{\end{theorem}}

\newcommand{\blemma}{\begin{lemma}}
\newcommand{\blemmarem}[1]{\begin{lemma}[#1]}
\newcommand{\elemma}{\end{lemma}}

\newcommand{\bexam}{\begin{examp}}
\newcommand{\eexam}{\end{examp}}

\newcommand\1{$\overline 1$}

\begin{document}

\title{On $Z_{2^k}$-Dual Binary Codes%
\thanks{This is author's version of the correspondence in the
IEEE Transactions on Information Theory 53(4) 2007, 1532--1537, Digital
Object Identifier \href{http://dx.doi.org/10.1109/TIT.2007.892787}{10.1109/TIT.2007.892787}, \copyright
2007 IEEE.}
\thanks{Some results of this paper were presented at
the 4th International Workshop on Optimal Codes and Related Topics
OC\,2005 (Pamporovo, Bulgaria, June 2005).}%
}
\renewcommand\footnotemark{ }
\author{Denis~S.~Krotov%
\thanks{D.\,S.\,Krotov is with the Sobolev Institute of Mathematics, Novosibirsk, Russia
(e-mail: {\tt krotov@math.nsc.ru}).}
}

\markboth
{Krotov: On $Z_{2^k}$-Dual Binary Codes}
{Krotov: On $Z_{2^k}$-Dual Binary Codes}

\maketitle

\begin{abstract}
A new generalization of the Gray map is introduced.
The new generalization $\Phi: Z_{2^k}^n \to Z_{2}^{2^{k-1}n}$
is connected with the known generalized Gray map $\varphi$
in the following way:
if we take two dual linear $Z_{2^k}$-codes and construct
binary codes from them using the generalizations $\varphi$ and $\Phi$ of the Gray map,
then the weight enumerators of the binary codes obtained will satisfy
the MacWilliams identity.
The classes of
$Z_{2^k}$-linear Hadamard codes and co-$Z_{2^k}$-linear
extended 1-perfect codes are described,
where co-$Z_{2^k}$-linearity means that the code can be obtained from
a linear $Z_{2^k}$-code with the help of the new generalized Gray map.
\begin{keywords}
Gray map, Hadamard codes, MacWilliams identity, perfect codes, $Z_{2^k}$-linearity
\end{keywords}
\end{abstract}

\section{Introduction}\label{S_Intro}
{A}{s} discovered in \cite{Nechaev82,Nechaev:Kerdock_cyclic_form}
certain nonlinear binary codes can be represented as linear codes over $Z_4$.
The variant of this representation founded in \cite{HammonsOth:Z4_linearity}
use the mapping $\phi: 0\to 00,\ 1\to 01,\ 2\to 11, 3\to 10$,
which is called the {\it Gray map},
to construct binary so-called $Z_4$-linear codes from linear quaternary codes.
The main property of $\phi$ from this point of view is that it is an isometry
between $Z_4$ with the Lee metric and $Z_2^2$ with the Hamming metric.
In \cite{Carlet:Z2k-linear} (and in \cite{HonNech} in more general form)
the Gray map is generalized to construct $Z_{2^k}$-linear codes.
The generalized Gray map
(say $\varphi$; see \hyperref[S_Z-lin0]{Subsection~}\ref{S_Z-lin0} for recalling basic facts
on the generalized Gray map)
is an isometric imbedding of $Z_{2^k}$ with the metric
specified by the homogeneous weight \cite{ConstHeise}
into $Z_2^{2^{k-1}}$ with the Hamming metric.

In this paper we introduce another generalization $\Phi$ of the Gray map
(\hyperref[S_co-Z-lin]{Subsection~}\ref{S_co-Z-lin}).
This generalization turns out to be dual to the previous in the following sense.
If $\cal C$ and ${\cal C}^\bot$ are dual linear $Z_{2^k}$-codes,
then the binary $Z_{2^k}$-linear
code $\varphi({\cal C})$ and the {\it co-$Z_{2^k}$-linear}
code $\Phi({\cal C}^\bot)$ are formally dual.
The formal duality is that the weight enumerators
of these two codes satisfy the MacWilliams identity (\hyperref[S_dual]{Section~}\ref{S_dual});
note that these codes, in general, can be nonlinear, in which case they cannot be
dual in the usual sense, as subspaces of the binary vector space.
So, we solve the problem of duality
for $Z_{2^k}$-linear binary codes: to relate the weight enumerators of the images
of dual linear codes over $Z_{2^k}$. This problem
cannot be solved using only the standard generalized Gray map $\varphi$,
because, as noted in \cite{Carlet:Z2k-linear},
the weight enumerators of $\varphi({\cal C})$ and $\varphi({\cal C}^\bot)$
are in general not related, contrarily to the case of $Z_4$-linear codes.

In \cite{HonNech}, it is shown that binary (and, in general, nonbinary)
codes can be represented as group codes over different groups.
Such representations use a special \emph{scaled isometry}
($\varphi$ is a partial case of such isometry),
which acts isometrically from
some module with a specially defined metric to the binary (or, in general, nonbinary)
Hamming space.
With such approach, each module element corresponds to one codeword;
$Z_{2^k}$-linear codes are a partial case of that approach.

In our approach, every module element (word) corresponds to some set of codewords,
which is a Cartesian product of the sets corresponding,
by a special mapping, to the symbols of the module word.
In the constructive part,
this approach can be represented
by the generalized concatenation construction \cite{Zin1976:GCC},
but the resulting code distance differs
from the constructive distance of generalized concatenated codes.
Calculating distance, we use some isometrical properties of the mapping, which
can be considered as an analog of a scaled isometry.
The map $\Phi$ considered in this paper allows to construct only codes with distance maximum $4$,
but in general the approach may have a larger potential.

The new way to generalize the Gray map (\hyperref[S_co-Z-lin]{Subsection~}\ref{S_co-Z-lin})
and its duality (\hyperref[S_dual]{Section~}\ref{S_dual}) to the ``old way''
are the main results of the first part of the paper.
The second part (\hyperref[S_perf]{Sections~}\ref{S_perf}\hyperref[S_Had]{--}\ref{S_nonexist})
is devoted to constructions of
co-$Z_{2^k}$-linear and $Z_{2^k}$-linear codes with parameters of
extended $1$-perfect and Hadamard codes.
There is a series
(e.\,g.
\cite{RifPuj:1997},
\cite{BorRif:1999},
\cite{Kro:2000:Z4_Perf},
\cite{Kro:2001:Z4_Had_Perf},
\cite{PheRif:2002},
\cite{BorPheRif:2003},
\cite{PheRifVil:2004ACCT:additive_hadamard},
\cite{PheRifVil:Hadamard-rank-kernel},
\cite{GuptaOth2005})
of papers that concern $1$-perfect (or extended $1$-perfect)
and Hadamard codes which are linear in some nonstandard sense
(note that for each considered values of parameters
only exists one linear code, up to equivalence).
Using the two generalized $\varphi$, $\Phi$ Gray maps, we construct a wide class of such codes.
Some natural questions on the constructed codes remains open for future researching:
which of these codes are equivalent to each other or to other known codes;
which of them are propelinear (see the definition in \cite{RifPuj:1997}) or
at least transitive
(see \cite{Sol:2005transitive} for the definition
and some constructions of transitive $1$-perfect codes);
to establish the bounds on the dimension of the kernel and rank
for these codes
(the dimension of the kernel and rank are good measures of linearity
of nonlinear codes in the classical binary sense.
For $Z_4$-linear extended $1$-per\-fect and Hadamard codes these parameters were calculated in
\cite{Kro:2000:Z4_Perf}, \cite{Kro:2001:Z4_Had_Perf}, \cite{BorPheRif:2003}, \cite{PheRifVil:2004ACCT:additive_hadamard}, \cite{PheRifVil:Hadamard-rank-kernel});
and so on.

In \hyperref[S_perf]{Section~}\ref{S_perf} we construct
co-$Z_{2^k}$-linear extended $1$-perfect $(n,2^n/2n,4)$ codes.
As noted above, the code distance of a co-$Z_{2^k}$-linear code
cannot exceed $4$ if $k>2$ (see \hyperref[DcoLin]{Lemma~}\ref{DcoLin}).
So, extended $1$-perfect codes are the best examples.

As a corollary, the dual distance of a $Z_{2^k}$-linear code ($k>2$) also
cannot exceed $4$. Best examples are codes with large code distance.
In \hyperref[S_Had]{Section~}\ref{S_Had} we construct $Z_{2^k}$-linear codes with
parameters $(n,2n,n/2)$, i.\,e., Hadamard codes.
They are
$Z_{2^k}$-dual to the extended $1$-perfect codes
from \hyperref[S_perf]{Section~}\ref{S_perf};
i.\,e., the $\varphi$ and, respectively, $\Phi$ preimage of the constructed
extended $1$-perfect and Hadamard codes are dual $Z_{2^k}$ codes.

The introduced constructions of codes are a generalization of
the constructions of $Z_4$-linear extended $1$-perfect and Hadamard codes
\cite{Kro:2000:Z4_Perf,Kro:2001:Z4_Had_Perf} (see also \cite{BorPheRif:2003}).

It is natural to ask whether our construction of
co-$Z_{2^k}$-linear extended $1$-perfect
and $Z_{2^k}$-linear Hadamard codes is complete or not.
In \hyperref[S_nonexist]{Section~}\ref{S_nonexist} we
show that the answer is yes;
i.\,e., all co-$Z_{2^k}$-linear extended $1$-perfect
codes and all $Z_{2^k}$-linear Hadamard codes are equivalent to codes constructed
in \hyperref[S_perf]{Sections~}\ref{S_perf} \hyperref[S_Had]{and~}\ref{S_Had}.

\section[Definitions And Basic Facts]{Definitions And Basic Facts. Two Generalizations Of Gray Map}\label{S_def}
{W}{e} will use the following common notations.
An \emph{$(n,M,d)$ binary code} is a cardinality $M$ subset of
$Z_2^n = {Z_2\times\ldots\times Z_2}$
with the distance at least
$d$ between every two different elements (\emph{codewords});
$n$ is called the code \emph{length} and $d$ is called the \emph{code distance}.
An $(n,2n,n/2)$ binary code is called an \emph{Hadamard code}; such codes exist
if and only if Hadamard $n\times n$ matrices exist, i.\,e.,
at least $n\equiv 0\mbox{ mod }4$ if $n>2$, see e.\,g. \cite{MWS}.
A linear Hadamard code is known as
a first order Reed-Muller code;
$n$ is a power of two is this case.
An $(n,2^n/2n,4)$ binary code is called an \emph{extended $1$-perfect code};
such codes exist if and only if $n$ is a power of two.
Such a code is always an extension of a $(n-1,2^n/2n,3)$ code, which is called
$1$-per\-fect; so, studying extended $1$-perfect codes is an alternative way
to study $1$-perfect codes. For each admissible length there is only one,
up to coordinate permutation, linear (extended) $1$-perfect code and,
if $n>8$, many nonlinear ones;
the classification of all such codes is currently  an open problem.
In the following two subsections
we will define the basic concepts, which will be used throughout the paper:
two weight functions on $Z_{2^k}$, two corresponding metrics,
two generalized Gray maps $\varphi$ and $\Phi$,
the concepts of $Z_{2^k}$-li\-ne\-ar and co-$Z_{2^k}$-li\-ne\-ar codes;
we will formulate simple but fundamental claims on isometric properties of $\varphi$ and $\Phi$.
\subsection[$Z_{2m}$-Linear Codes]{$Z_{2m}$-Linear Codes}\label{S_Z-lin0}

This subsection recalls basic definitions and facts about the
generalized Gray map and $Z_{2^k}$-linear codes \cite{Carlet:Z2k-linear,HonNech}.

Assume $m \geq 2$ is an integer such that there exist an Hadamard
$m\times m$ matrix.
Let $A\subset Z_2^{m}$ be a Hadamard $(m,2m,m/2)$ code.
Assume that $A=\{a_0,a_1, \ldots ,a_{2m-1}\}$, where $a_0$
is the all-zero codeword
and $a_i+a_{i+m}$ is the all-one codeword for each $i$ from $0$ to $m-1$.
Define the {\em generalized Gray map} $\varphi : Z_{2m}^n \to Z_{2}^{mn}$ by the rule
$$
\varphi(x_1,\ldots,x_n)\triangleq (a_{x_1},\ldots,a_{x_n}).
$$
Let the weight function $wt^*:Z_{2m}\to R^+$ be given by
$$
wt^*(x) \triangleq
\cases{
0& \mbox{if $x=0$}, \cr
m& \mbox{if $x=m$}, \cr
m/2& \mbox{otherwise}.}
$$
Whenever $m$ is a power of two, the weight $wt^*$ is the \emph{homogeneous weight} introduced in \cite{ConstHeise}
for more general class of rings. The corresponding distance $d^*$ on $Z_{2m}^n$
is defined by the standard way:
$$ d^*((x_1,\ldots,x_n),(y_1,\ldots,y_n))
\triangleq
\sum_{i=1}^n wt^*(y_i-x_i).
$$

\bprorem{\cite{HonNech}} \label{P_iso}
The mapping $\varphi$ is an isometric embedding of $(Z_{2m}^n,d^*)$ into
$(Z_2^{mn},d)$, i.\,e.,
$d^*(\bar x,\bar y)=d(\varphi(\bar x),\varphi(\bar y))$, where $d(\cdot,\cdot)$
is the Hamming distance.
\epro

If ${\cal C}\subseteq Z_{2m}^n$, then
$$
\varphi({\cal C})\triangleq \{ \varphi(\bar x) \,|\, \bar x\in {\cal C} \}.
$$

We will say that ${\cal C}$ is a $(n,M,d)^*$ code if 
${\cal C}\subseteq Z_{2m}^n$,
$|{\cal C}|=M$, and the $d^*$-distance between
any two different elements of ${\cal C}$ is not less than $d$.

\bcor\label{Cor1}
If ${\cal C}$ is a $(n,M,d)^*$ code, then $\varphi({\cal C})$
is a binary $(mn,M,d)$ code.
\ecor

A binary code is called {\em $Z_{2m}$-linear} if 
its coordinates can be arranged in such a way that
it is the image of a linear $Z_{2m}$-code by $\varphi$.
As noted in \cite{Carlet:Z2k-linear}, the length of a $Z_{2m}$-linear code
must be a multiple of $m$ and all the weights of its codewords
must be multiples of $m/2$.

\bnote
The map $\varphi$ and the concept
of $Z_{2m}$-linearity given above depends on the choice of
a Hadamard code $A$ and the only restriction on $m$ is that exists
a Hadamard $m\times m$ matrix.
In fact, originally {\rm \cite{Carlet:Z2k-linear}}
(and in the binary subcase of \cite[Section D]{HonNech})
$m=2^{k-1}$ and the code $A$ is fixed as
$R(1,k-1)$, the Reed-Muller code of order $1$.
\enote

\subsection[Co-$Z_{2^k}$-Linear Codes]{Co-$Z_{2^k}$-Linear Codes}
\label{S_co-Z-lin}
In this subsection we will introduce another approach
to construct binary codes from linear codes over $Z_{2^k}$.

Firstly, we will introduce a new way to generalize the Gray map, defining a map
$\Phi : Z_{2m}^n \to 2^{Z_{2}^{mn}}$,
where $2^{Z_{2}^{mn}}$ denotes the set of the subsets of $Z_{2}^{mn}$.
Then we will define a weight function $wt^\diamond:Z_{2m}\to R^+$
and the corresponding distance
$d^\diamond$ and we will show (\hyperref[DcoLin]{Lemma~}\ref{DcoLin})
relations between the $d^\diamond$-distance of a code ${\cal C}\subset Z_{2m}^n$
and the Hamming distance of the code $\Phi({\cal C})$.
Finally, we will introduce the concept of co-$Z_{2^k}$-linear codes.

Put $m=2^{k-1}$. Let $\{ H_0,\ldots,H_{2m-1} \}$ be a partition of $Z_2^{m}$ into
extended $1$-perfect $(m,2^m/2m,4)$ codes
(for example, we can take $H_0$ as the extended Hamming code and $H_0,\ldots,H_{2m-1}$ as its cosets).
Moreover, we assume that $H_0$ contains the all-zero word $\bar 0$ and
$H_j$ is an even (odd) weighted if and only if $j$ is even (odd).

Define the map $\Phi : Z_{2m}^n \to 2^{Z_{2}^{mn}}$ by the rule
$$
\Phi(x_1,\ldots,x_n)\triangleq H_{x_1}\times\ldots\times H_{x_n}.
$$
If ${\cal C}\subseteq Z_{2m}^n$, then\\[-10pt]
$$
\Phi({\cal C})\triangleq \bigcup_{\bar x\in {\cal C}} \Phi(\bar x).
$$

\bexam
Let $k=3$, $H_0=\{0000,1111\}$, \ldots,
$H_2=\{1100,0011\}$, \ldots,
 $H_7=\{0001,1110\}$. Then
\begin{IEEEeqnarray*}{rCrCl}
\Phi(2\,0\,7)&=&H_2\times H_0\times H_7
&=&\{\, 1100\,0000\,0001,\ 1100\,0000\,1110,\ 1100\,1111\,0001,\ 1100\,1111\,1110,\\
&&&&\phantom{\{\,} 0011\,0000\,0001,\ 0011\,0000\,1110,\ 0011\,1111\,0001,\ 0011\,1111\,1110\,\}.
\end{IEEEeqnarray*}
\eexam
\hspace{1.0ex}

Define the weight function $wt^\diamond:Z_{2m}\to R^+$ as
$$
wt^\diamond(x)\triangleq
\cases{
0&\mbox{if $x=0$},\cr
1&\mbox{if $x$ is odd},\cr
2&\mbox{if $x\neq 0$ is even}.}
$$
The corresponding distance $d^\diamond$ on $Z_{2m}^n$ is defined by the standard way:
$$ d^\diamond((x_1,\ldots,x_n),(y_1,\ldots,y_n))
\triangleq
\sum_{i=1}^n wt^\diamond(y_i-x_i).
$$

We will say that ${\cal C}$ is an $(n,M,d)^\diamond$ code if 
${\cal C}\subseteq Z_{2m}^n$,
$|{\cal C}|=M$, and the $d^\diamond$-distance between
any two different elements of ${\cal C}$ is not less than $d$.

\blemma\label{DcoLin}
Let $m\geq 4$.
If ${\cal C}\subseteq Z_{2m}^n$ is an $(n,M,d)^\diamond$ code,\\
then $\Phi({\cal C})$ is a binary $(mn,M (\frac{2^m}{2m})^n,\min(4,d))$ code.
\elemma
The proof is straightforward, and we omit it.
%
%
%
%
%
%

We call a binary code {\it co-$Z_{2m}$-linear} if 
its coordinates can be arranged so that
it is the image of a linear $Z_{2m}$-code by the map $\Phi$.

\bnote
Co-$Z_{2m}$-linear codes can be considered as a partial case
of generalized concatenated codes \cite{Zin1976:GCC}.
For this special case the code distance given by \hyperref[DcoLin]{Lemma~}\ref{DcoLin}
may be better than the code distance guaranteed by the
generalized concatenation construction.
\enote

\section[$Z_{2^k}$-Duality Of Binary Codes]{%
$Z_{2^k}$-Duality Of Binary Codes}
\label{S_dual}
{I}{n} this section we will show (\hyperref[duality]{Theorem~}\ref{duality})
that if two binary codes are obtained from
dual linear $Z_{2^k}$-codes by, both, $\varphi$ and $\Phi$
generalizations of Gray map,
then these codes are formally dual, i.\,e., their weight enumerators satisfy the
MacWilliams identity.

Let $C$ be a $Z_{2^k}$-linear code, the image by $\varphi$
of a linear $Z_{2^k}$-code $\cal C$ of length $n$.
Let ${\cal C}^\bot$ be the linear $Z_{2^k}$-code dual to ${\cal C}$.
Let $SW_{{\cal C}^\bot}(X,Z,T)$ be the polynomial obtained from the complete
weight enumerator $W_{{\cal C}^\bot}(X_0,X_1,\ldots,X_{2^k-1})$ of ${\cal C}^\bot$
by identifying to $Z$ (respectively to $T$)
all the $X_j$'s such that $j$ is odd (respectively, $j$ is even $\neq 0$).

\blemmarem{\cite{Carlet:Z2k-linear}}\label{CC}
Then we have
\begin{IEEEeqnarray*}{rRl}
W_C(X,Y)&
=&\frac{1}{|{\cal C}^\bot|}SW_{{\cal C}^\bot} \big(\
X^{2^{k-1}}\hspace{-0.8ex}+Y^{2^{k-1}}\hspace{-0.8ex}+(2^k-2)(X Y)^{2^{k-2}}\hspace{-0.4ex},\ \nonumber \\
&&\phantom{\frac{1}{|{\cal C}^\bot|}SW_{{\cal C}^\bot} \big(\ }
X^{2^{k-1}}\hspace{-0.8ex}-Y^{2^{k-1}},\  \nonumber \\
&&\phantom{\frac{1}{|{\cal C}^\bot|}SW_{{\cal C}^\bot} \big(\ }
X^{2^{k-1}}\hspace{-0.8ex}+Y^{2^{k-1}}\hspace{-0.8ex}-2(X Y)^{2^{k-2}}\ \big).\nonumber \\
\end{IEEEeqnarray*}
\elemma

The following lemma is a known fact about the weight distribution of
extended $1$-perfect binary codes.
\blemma\label{W_H}
Let $H$ be an extended $1$-perfect $(m,2^m/2m,4)$ code;
then \\
{\rm a)}
\(
\frac{1}{|H|}
W_H(X+Y,X-Y)=X^{m}+Y^{m}+(2m-2)(X Y)^{m/2}
\) \ if $\bar 0\in H$,
\\[0pt]
{\rm b)}
\(
\frac{1}{|H|}
W_H(X+Y,X-Y)=X^{m}-Y^{m}
\) \ if $H$ is odd-weight,
\\[0pt]
{\rm c)}
\(
\frac{1}{|H|}
W_H(X+Y,X-Y)=X^{m}+Y^{m}-2(X Y)^{m/2}
\) \ if $H$ is an even-weight code and $\bar 0\not\in H$.%
\elemma

\bpro\label{W-SW}
Let $\cal C$ be a $Z_{2^k}$-code and
$\tilde C=\Phi({\cal C})$; then
$$
W_{\tilde C}(X,Y)=SW_{\cal C}(W_{H_0}(X,Y), W_{H_1}(X,Y), W_{H_2}(X,Y)),
$$
where $\{ H_0,\ldots,H_{2m-1} \}$ is the partition of $Z_2^{m}$ into
extended $1$-perfect codes from the definition of $\Phi$,
$0\in H_0$, $H_1$ is an odd-weight code,
$H_2$ is an even-weight code, $\bar 0\not\in H_2$
\epro
\proofr
The statement follows almost immediately from the definition of the map $\Phi$.
Indeed, each codeword $\bar z=(z_1,\ldots,z_n)\in {\cal C}$ adds
$SW_{z_1}\cdot\ldots\cdot SW_{z_n}$ to $SW_{\cal C}(X,Z,T)$, where
\begin{IEEEeqnarray*}{rCl"l}
  SW_{0} &\triangleq& X, \\
  SW_{2j+1}  &\triangleq& Z,& j=0, \ldots ,m-1,\\
  SW_{2j}  &\triangleq& T, & j=1, \ldots ,m-1.
\end{IEEEeqnarray*}
On the other hand, as follows from the definition of the map $\Phi$,
each codeword $\bar z=(z_1,\ldots,z_n)\in {\cal C}$ adds
$W_{H_{z_1}}(X,Y)\cdot\ldots\cdot W_{H_{z_n}}(X,Y)$ to $W_{\tilde C}(X,Y)$.
The relations
\begin{IEEEeqnarray*}{rCl"l}
  W_{H_{2j+1}}(X,Y)&=&W_{H_{1}}(X,Y), \qquad j=0, \ldots ,m-1,\\
  W_{H_{2j}}(X,Y)&=&W_{H_{2}}(X,Y), \qquad j=1, \ldots ,m-1
\end{IEEEeqnarray*}
follow from \hyperref[W_H]{Lemma~}\ref{W_H} and conclude the proof.
\proofend

The following theorem is the main result of this section.
\btheorem\label{duality}
Let ${\cal C}$ and ${\cal C}^\bot$ be dual linear $Z_{2^k}$-codes,
$C=\varphi({\cal C})$ and $\tilde C_\bot=\Phi({\cal C}^\bot)$.
Then the codes $C$ and $\tilde C_\bot$ are formally dual, i.\,e.,
\(
 W_C(X,Y)=\frac{1}{|\tilde C_\bot|}W_{\tilde C_\bot}(X+Y,X-Y).
\)
\etheorem
\proofr
By \hyperref[CC]{Lemmas~}\ref{CC} \hyperref[W_H]{and~}\ref{W_H}
\begin{IEEEeqnarray*}{rCl}
W_C(X,Y)
&=&\frac{1}{|{\cal C}^\bot|}SW_{{\cal C}^\bot} \bigg(
\frac{1}{|H_0|}W_{H_0}(X{+}Y,X{-}Y),
\frac{1}{|H_1|}W_{H_1}(X{+}Y,X{-}Y),
\frac{1}{|H_2|}W_{H_2}(X{+}Y,X{-}Y)
\bigg)\\
&=&\frac{1}{|{\cal C}^\bot|}\left(\frac{2m}{2^m}\right)^{\!\!n}SW_{{\cal C}^\bot} \Big(
W_{H_0}(X{+}Y,X{-}Y),
W_{H_1}(X{+}Y,X{-}Y),
W_{H_2}(X{+}Y,X{-}Y)
\Big).
\end{IEEEeqnarray*}
It remains to note that by \hyperref[DcoLin]{Lemma~}\ref{DcoLin} we have
${|{\cal C}^\bot|}\left(\frac{2^m}{2m}\right)^n =|\tilde C_\bot|$
and by \hyperref[W-SW]{Proposition~}\ref{W-SW} we obtain:
\begin{IEEEeqnarray*}{rlCl}
SW_{{\cal C}^\bot} \big(W_{H_0}(X{+}Y,X{-}Y),
W_{H_1}(X{+}Y,X{-}Y),
W_{H_2}(X{+}Y,X{-}Y){\big)}
{}={}&W_{\tilde C_\bot}(X{+}Y,X{-}Y).
\vspace{-4.1ex} 
\end{IEEEeqnarray*}
\vspace{2ex} 
\proofend
In \hyperref[S_perf]{Sections~}\ref{S_perf} \hyperref[S_Had]{and~}\ref{S_Had}
we will construct two $Z_{2^k}$-dual classes of codes:
co-$Z_{2^k}$-linear extended $1$-perfect codes
and $Z_{2^k}$-linear Hadamard codes.
In the \hyperref[S_nonexist]{last section}
we will show the completeness of the constructions.
\section[Co-$Z_{2^k}$-Linear Extended $1$-Perfect Codes]{%
Co-$Z_{2^k}$-Linear Extended Perfect Codes}
\label{S_perf}

{T}{his} section concerns extended $1$-perfect codes.
We first introduce the concept of {\em \1-perfect code}, which is a generalization of
the concept of extended $1$-perfect binary code to some nonbinary cases.
As in the case of $1$-per\-fect codes, the existence of \1-per\-fect codes
in different spaces is independently interesting.
Then, in \hyperref[ss_b]{Subsection~}\ref{ss_b},
we  construct a class of {\em \1-perfect codes} in $(Z_{2^k}^n,d^\diamond)$.
In \hyperref[ss_c]{Subsection~}\ref{ss_c} we  summarize:
the images of such codes under $\Phi$ are co-$Z_{2^k}$-linear extended $1$-perfect codes.
In \hyperref[ss_d]{Subsection~}\ref{ss_d}
we give examples of \1-per\-fect codes in $Z_{2m}$ where $m$ is not a power of two.

\subsection[$1'$-Perfect Codes]{\1-Perfect Codes} \label{ss_a}
Let $G=(V,E)$ be a regular bipartite graph with parts $V_{ev}$, $V_{od}$.
A subset $C\subseteq V_{ev}$ is called a {\em \1-perfect code} if 
for each $\bar x\in V_{od}$ there exist exactly one $\bar c\in C$
adjacent with $\bar x$.
It is not difficult to see that a \1-perfect code is an optimal distance $4$ code,
i.\,e., its cardinality is maximum among the distance $4$ codes in the same space.
In particular, such a code is a $3$-diameter perfect code in the sense \cite{AhlAydKha}.
\1-perfect codes in a binary Hamming space are known as extended $1$-perfect codes.
The following test can be considered as an alternative definition of a \1-perfect code.
\bpro\label{1perf}
Assume $G=(V,E)$ is a regular bipartite graph  of degree $r>0$
and $d^{G}$ is the graph distance in $V$.
Then $C\subset V$ is a \1-perfect code if and only if
$|C|=|V|/2r$ and $d^{G}(\bar c_1,\bar c_2)\geq 4$
for each different $\bar c_1, \bar c_2\in C$.
\epro
\proofr
{\em Only if:} Assume $C\subset V$ is a \1-perfect code.
Then by the definition it is a subset of a part $V_{ev}$
of the graph $(V,E)$ and the graph distance between any two elements of $C$ is even.

On the other hand, if $d^{G}(\bar c_1, \bar c_2)=2$ for some $\bar c_1, \bar c_2\in C$,
then there exist an element $\bar x\in V$ such that
$d^{G}(\bar x, \bar c_1)=1$ and $d^{G}(\bar x, \bar c_2)=1$.
The existence of such element contradicts to the definition of \1-perfect code.
Consequently, $d^{G}(\bar c_1, \bar c_2)\geq 4$ for each different $\bar c_1, \bar c_2\in C$.

Each element of $V_{od}\triangleq V\backslash V_{ev}$ is adjacent with exactly one element of $C$.
On the other hand, each element of $C$ is adjacent with exactly $r$ elements of $V_{od}$.
Therefore, $|C|=|V_{od}|/r=|V|/2r$.

{\em If:}
Assume $C$ is a code of cardinality $|V|/2r$ and distance at least $4$.
Denote
$$
C_1 \triangleq \{\bar x\,|\,d^{G}(\bar x,C)=1\}.
$$
Each element of $C_1$ is adjacent with exactly one element of $C$
(otherwise the code distance of $C$ is not more than $2$).
Consequently, $|C_1|=r|C|=|V|/2$.

(*) {\em We claim that $C_1$ does not contain two adjacent elements.}
Assume the contrary.
Then the graph $G$ contains a chain $(\bar c, \bar x, \bar y, \bar c')$,
where $\bar c, \bar c'\in C$, $\bar x, \bar y \in C_1$.
The case $\bar c=\bar c'$ contradicts to the bipartiteness of the graph $G$.
The case $\bar c\neq\bar c'$ contradicts to the code distance of $C$.
The claim (*) is proved.

Since $G$ is a regular graph, the other half of vertices $V \backslash C_1$
also does not contain two adjacent elements.
Therefore $C$ is a \1-perfect code by the definition, where $V_{od}=C_1$.
\proofend

\subsection[A Class Of $1'$-Perfect Codes In $Z_{2^k}^n$]{A Class Of \1-Perfect Codes In $Z_{2^k}^n$} \label{ss_b}
Let $n=2^r$ and $I=(i_1,\ldots,i_k)$ be a collection of nonnegative integers such that
$1 i_1+2 i_2+ \ldots +k i_k=r$.
Let $\bar b_1,\ldots,\bar b_n\in Z_{2^k}^{1+i_1+...+i_k}$ be all the elements
of $\{1\}\times (2^{k-1} Z_{2^k})^{i_1}\times (2^{k-2} Z_{2^k})^{i_2}
\times \ldots \times (2^{0} Z_{2^k})^{i_k}$ ordered lexicographically.
And let $B_I$ be the matrix with the columns $\bar b_1,\ldots,\bar b_n$.

\bexam
If $k=3$, $i_1=2$, $i_2=1$, $i_3=0$, then $r=4$, $n=2^r=16$, and
$$
B_I=
\left(%
\begin{array}{c}
1111111111111111 \\
0000000044444444 \\
0000444400004444 \\
0246024602460246 \\
\end{array}
\right).
$$
If $k=3$, $i_1=0$, $i_2=i_3=1$, then $r=5$, $n=2^r=32$, and
$$
B_I=
\left(%
\begin{array}{c}
11111111111111111111111111111111 \\
00000000222222224444444466666666 \\
01234567012345670123456701234567 \\
\end{array}
\right).
$$
If $k=3$, $i_1=i_2=i_3=1$, then $r=6$, $n=2^r=64$, and
$$
B_I=
\left(%
\begin{array}{c}
1111111111111111111111111111111111111111111111111111111111111111 \\
0000000000000000000000000000000044444444444444444444444444444444 \\
0000000022222222444444446666666600000000222222224444444466666666 \\
0123456701234567012345670123456701234567012345670123456701234567 \\
\end{array}
\right).
$$
\eexam

\blemma\label{isPerf}
The linear code
${\cal H}_I\triangleq \{\bar h=(h_1,\ldots,h_n)\in Z_{2^k}^n\,
|\,\sum_{j=1}^n h_j \bar b_j = B_I\bar h^T = \bar 0 \}$ with the check matrix $B_I$
is \1-perfect.
\elemma
\proofr
First we claim that

(*) {\em the distance $2$ between codewords
$\bar h_1$, $\bar h_2$ from ${\cal H}_I$ is impossible}.
Indeed, if $d^\diamond(\bar h_1,\bar h_2)=2$, then the codeword
$\bar h \triangleq \bar h_1-\bar h_2$ have one or two nonzero positions.
The first case contradicts to the fact that $\bar b_i\neq \bar 0$, $i=1, \ldots ,n$.
In the second case we have $\beta \bar b_i+\gamma \bar b_j=\bar 0$ for some different $i$, $j$
and odd $\beta$, $\gamma$. But the first row of $B_I$ implies that $\beta=-\gamma$ and,
since $\beta$ and $\gamma$ are odd, we get $\bar b_i=\bar b_j$.
This again contradicts to the construction of $B_I$. The claim (*) is proved.

First row of 
$B_I$ implies that the words of ${\cal H}_I$ have
even weight.

We need to check that each
odd word is at the distance $1$ from exactly one codeword.
Let $\bar z=(z_1, \ldots ,z_n)\in Z_{2^k}^n$ have an odd weight,
$\bar s\triangleq\sum_{j=1}^n z_i\bar b_i$, and $s\triangleq\sum_{j=1}^n z_i$.
Note that $s$ is odd.
Since $s^{-1}\bar s\in \{1\}\times (2^{k-1} Z_{2^k})^{i_1}\times (2^{k-2} Z_{2^k})^{i_2}
\times \ldots \times (2^{0} Z_{2^k})^{i_k}$, there exists $j'$ such that
$s^{-1}\bar s=\bar b_{j'}$. Let $\bar z'\in Z_{2^k}^n$ be the word with $s$ in the $j'$th position
and zeroes in the others.
It is easy to check that $\bar z-\bar z'$ is a codeword at the distance $1$ from $\bar z$.
As follows from (*), such codeword is unique for each odd $\bar z$.
\proofend

\subsection{Co-$Z_{2^k}$-Linear Extended $1$-Perfect Binary Codes} \label{ss_c}

Now we have all we need to construct a class of co-$Z_{2^k}$-linear
extended $1$-perfect binary codes.
As we will see in \hyperref[S_nonexist]{Section~}\ref{S_nonexist},
the class constructed exhaust all such codes provided the mapping $\Phi$ is fixed.

\btheorem\label{perf}
The co-$Z_{2^k}$-linear code $\tilde H_I\triangleq \Phi({\cal H}_I)$ is
a binary $(nm,2^{nm}/2nm,4)$ code, i.\,e., an extended $1$-perfect code.
\etheorem

\proofr
The statement follows directly from
\hyperref[DcoLin]{Lem\-ma~}\ref{DcoLin},
\hyperref[DcoLin]{Pro\-po\-si\-ti\-on~}\ref{1perf}, and
\hyperref[DcoLin]{Lem\-ma~}\ref{isPerf}.
\proofend

\subsection{Note On The General Case $Z_{2m}$} \label{ss_d}

Indeed, \1-perfect codes can be constructed
over $Z_{2m}$ with $d^\diamond$-distance
for each $m\geq 1$, see the following examples.
Since $m=2^\mu$ is a necessary condition for constructing binary codes
using the way of \hyperref[S_co-Z-lin]{Subsection~}\ref{S_co-Z-lin},
the classification of \1-perfect codes
in the other cases is out of view of this paper.

\bexam\label{ex1perf} The codes with the check matrices

$$B'\triangleq
\left(%
\begin{array}{cccccccc}
  1 & 1 & 1 & 1 & 1 & 1 & 1 & 1  \\
  0 & 0 & 0 & 0 & 12& 12& 12& 12 \\
  0 & 6 & 12& 18& 0 & 6 & 12& 18
\end{array}%
\right),
$$
$$
B''\triangleq
\left(%
\begin{array}{cccccc}
  1 & 1 & 1  & 1  & 1  & 1  \\
  0 & 4 & 8  & 12 & 16 & 20
\end{array}%
\right)
$$
are \1-perfect codes over $Z_{24}$ with $d^\diamond$-distance.
\eexam

\section[$Z_{2^k}$-Linear {H}a\-da\-mard Codes]{%
$Z_{2^k}$-Linear {H}a\-da\-mard Codes}\label{S_Had}

\blemma \label{L_HD}
Let ${\cal H}$ be a linear code in $Z_{2^k}^n$.
Then ${\cal H}$ is an $(n,n2^{k},n2^{k-2})^*$ code
if and only if ${\cal H}^\bot$ is
a \1-perfect code in $(Z_{2^k}^n,d^\diamond)$.
\elemma
\proofr
Assume the code ${\cal H}$ has parameters $(n,n2^{k},n2^{k-2})^*$.
Then $\varphi({\cal H})$ has the parameters of an Hadamard code,
see \hyperref[Cor1]{Corollary~}\ref{Cor1}. By \hyperref[duality]{Theorem~}\ref{duality}
the code $\Phi({\cal H}^\bot)$ is formally dual to the Hadamard code $\varphi({\cal H})$, i.\,e,
$\Phi({\cal H}^\bot)$ is an extended $1$-perfect code. Then the cardinality and
the minimal $d^\diamond$-distance $4$ of the code ${\cal H}^\bot$
follow from \hyperref[DcoLin]{Lemma~}\ref{DcoLin}.

The reverse statement can be proved by reversing the arguments.
\proofend

The next theorem follows immediately from \hyperref[isPerf]{Lemmas~}\ref{isPerf}
\hyperref[L_HD]{and~}\ref{L_HD}.

\btheorem
The code
${\cal D}_I\triangleq {\cal H}_I^\bot$
is a linear $(n,n2^{k},n2^{k-2})^*$ code with the generator matrix $B_I$.
The $Z_{2^k}$-linear code $D_I\triangleq \varphi({\cal D}_I)$ is
a binary $(n2^{k-1},n2^{k},n2^{k-2})$ code, i.\,e., an Ha\-da\-mard code.
\etheorem

\bnote
The code ${\cal D}_{(r,0,\ldots,0)}$ is the
{\em first order Reed-Muller code over $Z_{2^k}$} \cite{GuptaOth2005}.
\enote

\bnote Indeed, if an Hadamard code
$A$ of length $m$ exists (see \hyperref[S_Z-lin0]{Subsection~}\ref{S_Z-lin0}),
then $Z_{2m}$-linear Hadamard code can be constructed for length $nm=2^r m$,
even if $m$ is not a power of two.
For example, the linear $Z_{24}$-code with the generator matrix $B'$
from  \hyperref[ex1perf]{Example~}\ref{ex1perf}
has parameters $(16,192,48)^*$ and the corresponding binary $Z_{24}$-linear
code has parameters $(96,192,48)$, i.\,e., the parameters of an Hadamard code of length $96$.

However, the code with the generator matrix $B''$ (\hyperref[ex1perf]{Example~}\ref{ex1perf})
has parameters $(6,144,30)^*$, corresponding a binary $(72,144,30)$ code.
The code distance of this code is smaller than the code distance of an Hadamard code
with the same length and cardinality.
\enote

\section[Nonexistence Of Unknown Co-$Z_{2^k}$-Linear Extended Perfect Codes And
$Z_{2^k}$-Linear {H}a\-da\-mard Codes]{%
Nonexistence Of Unknown Co-$Z_{2^k}$-Linear Extended Perfect Codes And
$Z_{2^k}$-Linear {H}a\-da\-mard Codes}\label{S_nonexist}
{L}{et} $n$ be a power of $2$. In this section we will show
(\hyperref[Th_HH]{Theorem~}\ref{Th_HH}) that
each linear $(n, 2^{kn}/n2^k, 4)^\diamond$ code in $Z_{2^k}^n$ is equivalent to a code
from the class constructed in \hyperref[S_perf]{Section~}\ref{S_perf}. Similarly,
each linear
$(n,n2^{k},n2^{k-2})^*$
 code in $Z_{2^k}^n$ is equivalent to a code
from the class constructed in \hyperref[S_Had]{Section~}\ref{S_Had}
(\hyperref[Th_DD]{Theorem~}\ref{Th_DD}). The key moment of the proof
is \hyperref[Th_HH]{Lemma~}\ref{L_1111}.
The partial $Z_4$ case of this lemma
was proved in \cite{Kro:2000:Z4_Perf} and in \cite{BorPheRif:2003},
but the proof given here is not similar.

We say that two linear codes ${\cal C}_1,{\cal C}_2\subseteq Z_{2^k}^n$
are \emph{equivalent} if 
${\cal C}_2=\bar z\circ\pi {\cal C}_1$
where $\pi$ is a coordinate permutation,
$\bar z \in (Z_{2^k}^*)^n\triangleq \{1,3,\ldots,2^k-1\}^n$,
and $\circ$ is the coordinate-wise product defined as
$(z_1,\ldots,z_n)\circ(x_1,\ldots,x_n)\triangleq (z_1 x_1,\ldots,z_n x_n)$.
Note that both operations $\pi$ and $\bar z\,\circ$ are group automorphisms of the
additive group of $Z_{2^k}^n$ and isometries of the metric spaces
$(Z_{2^k}^n,d^*)$ and $(Z_{2^k}^n,d^\diamond)$.
The following proposition shows that there are no other linear isometries
of $(Z_{2^k}^n,d^*)$ or $(Z_{2^k}^n,d^\diamond)$, and thus our definition
of the equivalence is natural.
\bpro
Assume that $\Gamma$ is a linear transformation of $Z_{2^k}^n$ and
isometry of $(Z_{2^k}^n,d)$ where $d=d^*$ or $d=d^\diamond$.
Then $\Gamma(\bar x)\equiv\bar z\circ\pi(\bar x)$
where $\pi$ is a coordinate permutation,
and $\bar z \in (Z_{2^k}^*)^n$.
\epro
\proofr
Denote by $\bar e_i$ the word with $1$ in $i$th
position and zeroes in the other positions.
It is enough to check that for each $i$ we have
$\Gamma(\bar e_i)\equiv z_j \bar e_{j}$ with some
$j=\pi(i)$ and $z_{j}\in Z_{2^k}^*$.
Indeed, from the isometric properties of $\Gamma$
we derive that $\Gamma(\bar e_i)$ has only one non-zero
coordinate.
On the other hand, this coordinate belongs to $Z_{2^k}^*$,
because $\Gamma$ is a bijection.
\proofend

The following two theorems are the main results of this section.
Remind that the codes ${\cal H}_I$ and ${\cal D}_I$ are defined in
\hyperref[S_perf]{Sections~}\ref{S_perf} \hyperref[S_Had]{and~}\ref{S_Had}.

\btheorem \label{Th_HH}
Let ${\cal H}\subset Z_{2^k}^n$ be a linear
\1-perfect
code.
Then ${\cal H}$ is equivalent to ${\cal H}_I$ where
$I=(i_1,\ldots,i_k)$ is a collection of nonnegative integers such that
$1 i_1+2 i_2+ \ldots +k i_k=\log_2 n$.
\etheorem

\btheorem \label{Th_DD}
Let ${\cal D}\subset Z_{2^k}^n$ be a linear $(n,n2^{k},n2^{k-2})^*$ code.
Then ${\cal D}$ is equivalent to ${\cal D}_I$, where
$I=(i_1,\ldots,i_k)$ is a collection of nonnegative integers such that
$1 i_1+2 i_2+ \ldots +k i_k=\log_2 n$.
\etheorem

By \hyperref[L_HD]{Lemma~}\ref{L_HD}
it is enough to prove only one of \hyperref[Th_HH]{Theorems~}\ref{Th_HH}, \ref{Th_DD}.
We need the following auxiliary result.

\blemma\label{L_1111}
If $\cal D$ is a linear $(n,n2^{k},n2^{k-2})^*$ code in $Z_{2^k}^n$, then
$\cal D$ contains an element from $(Z_{2^k}^*)^n$.
\elemma
\proofr
We prove the lemma by induction.
If $n=1$,
then ${\cal D}=Z_{2^k}$.
Assume $n>1$.

(*) \emph{We claim that $\cal D$ contains an element from $\{0,2^{k-1}\}^n$ of weight $n2^{k-2}$.}
Let $\bar x'$ and $\bar x''$ be two linear independent elements in $\cal D$
of order $2m'$ and $2m''$ respectively.
This means that $m' \bar x'$ and $m'' \bar x''$
are different nonzero elements from ${\cal D} \cap \{0,2^{k-1}\}^n$.
Since $\varphi$ is an isometric imbedding (see \hyperref[P_iso]{Proposition~}\ref{P_iso})
and $\varphi({\cal D})$ is an $(n2^{k-1},n2^{k},n2^{k-2})$ Hadamard code,
the only possible values of
$wt^*$-weight of elements in $\cal D$ are $0$, $n2^{k-2}$ and $n2^{k-1}$.
So, at least one of $m' \bar x'$ and $m'' \bar x''$ has weight $n2^{k-2}$.
The claim (*) is proved.

Without loss of generality
assume that
\(\bar a \triangleq (2^{k-1},\ldots,2^{k-1},0,\ldots,0)
\in {\cal D}.\)

Let us consider two codes obtained from $\cal D$ by puncturing $n/2$ coordinates:
\begin{IEEEeqnarray*}{rCl}
{\cal D}_1 & {}\triangleq{} & \{\bar z'\in Z_{2^k}^{n/2} \, | \  %
\exists\, \bar z''\in Z_{2^k}^{n/2} :\, (\bar z',\bar z'')\in {\cal D}\}, \\
{\cal D}_2 & {}\triangleq{} & \{\bar z''\in Z_{2^k}^{n/2} \, | \  %
\exists\, \bar z'\in Z_{2^k}^{n/2} :\, (\bar z',\bar z'')\in {\cal D}\}.
\end{IEEEeqnarray*}

(**) \emph{We claim that ${\cal D}_1$ and ${\cal D}_2$ are linear $(n/2,n2^{k-1},n2^{k-3})^*$ codes.}
The linearity of the codes is obvious.
The code distance follows from the fact that the distance between $\bar a$
and any element $\bar z\in {\cal D}$ is $0$, $n2^{k-2}$ or $n2^{k-1}$.
It is true that $|{\cal D}_1| \geq |{\cal D}|/2$ because the code distance of ${\cal D}$
permits of only one nonzero codeword with zeroes in the first $n/2$ coordinates.
On the other hand, such a codeword exists:
$(0,\ldots,0,2^{k-1},\ldots,2^{k-1})=(2^{k-1},\ldots,2^{k-1})+\bar a\in {\cal D}$.
So, $|{\cal D}_1| = |{\cal D}|/2 = n2^{k-1}$.
Similarly, we get $|{\cal D}_2| = n2^{k-1}$. The claim (**) is proved.

By the assumption of induction ${\cal D}_1$ contains an element $\bar u'$
from $(Z_{2^k}^*)^{n/2}$.
This means that there exist $\bar u''$ from $Z_{2^k}^{n/2}$ such that
$(\bar u',\bar u'')\in {\cal D}$.
Since $wt^*(2^{k-1}(\bar u',\bar u''))\in \{n2^{k-2},n2^{k-1}\}$ and
$2^{k-1}\bar u'=(2^{k-1},...,2^{k-1})$,
we have $wt^*(2^{k-1}\bar u'')=0$ or $wt^*(2^{k-1}\bar u'')=n2^{k-2}$.
So,
\setlength{\arraycolsep}{0mm}
\begin{eqnarray}
 \label{u''0} 2^{k-1}\bar u'' & {}={} & (0,\ldots,0), \\
 \label{u''1} \mbox{or}\ \ 2^{k-1}\bar u'' & {}={} & (2^{k-1},\ldots,2^{k-1}).
\end{eqnarray}
Similarly, considering the code ${\cal D}_2$, we can
find a codeword $(\bar v',\bar v'')\in {\cal D}$ such that
$2^{k-1}\bar v''=(2^{k-1},...,2^{k-1})$ and $v'$ satisfies\\[-20pt]
\setlength{\arraycolsep}{0mm}
\begin{eqnarray}
 \label{v'0} 2^{k-1}\bar v' & {}={} & (0,\ldots,0), \\
 \label{v'1} \mbox{or}\ \ 2^{k-1}\bar v' & {}={} & (2^{k-1},\ldots,2^{k-1}).
\end{eqnarray}
If (\ref{u''1}) is true, then $(u',u'')\in (Z_{2^k}^*)^{n}\cap {\cal D}$.
If (\ref{v'1}) is true, then $(v',v'')\in (Z_{2^k}^*)^{n}\cap {\cal D}$.
If (\ref{u''0}) is true and (\ref{v'0}) is true, then
$(u',u'')+(v',v'')\in (Z_{2^k}^*)^{n}\cap {\cal D}$.
\hyperref[L_1111]{Lemma~}\ref{L_1111} is proved.
\proofend

\proofrem{Proof of \hyperref[Th_DD]{Theorem~}\ref{Th_DD}}
By \hyperref[L_1111]{Lemma~}\ref{L_1111} the code
$\cal D$ contains an element $\bar c=(c_1,\ldots,c_n)$ from $(Z_{2^k}^*)^n$.
Then the code ${\cal D}'\triangleq (c_1^{-1},\ldots,c_n^{-1})\circ {\cal D}$
is equivalent to ${\cal D}$ and contains $\bar 1=(1,\ldots,1)$.
Let $\{\bar 1, q_1,\ldots, q_s\}$ be a basis of ${\cal D}'$ and
$i_j$ be the number of elements of order $2^{j}$, $j=1,\ldots,k$.
Assume $\bar 1, q_1,\ldots, q_s$ are the rows of the matrix $Q$
of size $(1+i_1+\ldots+i_k)\times n$, a generator matrix of the code ${\cal D}'$.
Then the columns of $Q$ are elements
of $\{1\}\times (2^{k-1} Z_{2^k})^{i_1}\times (2^{k-2} Z_{2^k})^{i_2}
\times \ldots \times (2^{0} Z_{2^k})^{i_k}$.
Since by \hyperref[L_HD]{Lemma~}\ref{L_HD} the code $d^\diamond$-distance of ${\cal D}'^\bot$ is more than $2$,
all the columns are pairwise different.
Therefore $Q$ coincides with $B_I$, $I=(i_1,\ldots,i_k)$,
up to permutation of columns.
\proofremend

So, we conclude, provided the mappings $\varphi$ and $\Phi$ are fixed,
all up to equivalence
co-$Z_{2^k}$-linear extended $1$-perfect codes and
$Z_{2^k}$-linear Hadamard codes are described
in \hyperref[S_perf]{Sections~}\ref{S_perf} \hyperref[S_Had]{and~}\ref{S_Had}.

\ifx\href\undefined \newcommand\href[2]{#2} \fi\ifx\url\undefined
  \newcommand\url[1]{\href{#1}{#1}} \fi

\end{document}